# SciPostLayoutTree: A Dataset for Structural Analysis of Scientific Posters

### v1

@@ 3.2 Dataset Annotation
We construct SciPostLayoutTree by adding structural annotations to SciPostLayout [26]. In contrast to business or artistic posters, scientific posters have an explicit and consistent reading order intended by the authors. This property allows us to define a unique DFS-ordered tree structure for each poster. As annotation errors can be objectively detected, inter-annotator agreement is not critical. Each poster was annotated as a DFS-ordered tree rooted at the Root node [^1] See the Appendix for the annotation guideline.. The annotation was conducted by professional annotators from an external vendor, with one annotator assigned per poster. All annotations were reviewed by in-house supervisors and the authors with expertise in scientific poster layouts, ensuring rigorous quality control. Posters containing only a single BBox were excluded as structurally trivial. As a result, we obtained 7,851 annotated posters, split into 6,853/499/499 for training, validation, and test sets. We confirmed that 98% of posters in the validation and test sets have non-overlapping authors. The test set size of 499 posters is sufficient compared to prior datasets such as DocHieNet.
@@ 3.3.1 Layout Element Statistics
Title 7,844 1.00 ($\pm$ 0.04)
Author Info 7,694 0.98 ($\pm$ 0.15)
Section 41,344 5.27 ($\pm$ 2.44)
Text 52,710 6.71 ($\pm$ 4.88)
List 23,760 3.03 ($\pm$ 3.31)
Figure 38,642 4.92 ($\pm$ 3.62)
Table 5,657 0.72 ($\pm$ 1.26)
Caption 14,974 1.91 ($\pm$ 2.67)
All 192,625 24.54 ($\pm$ 9.58)
@@ 5.1 Experimental Settings
We compared four DRGG variants, each with a different combination of the BBox Embedding and beam search.
@@ 5.2 Evaluation Metrics
ResNet-50 DRGG 68.75 75.05 8.82
DRGG-BE 84.46 86.61 4.31
DRGG-BS 76.83 83.12 6.66
DRGG-BEBS 88.45 90.38 3.20
ViT DRGG 80.41 85.84 5.41
DRGG-BE 86.24 88.11 3.86
DRGG-BS 83.78 89.25 4.49
DRGG-BEBS 89.67 91.34 2.86
Swin DRGG 79.84 85.78 5.70
DRGG-BE 87.14 88.86 3.57
DRGG-BS 83.11 89.02 4.75
DRGG-BEBS 89.50 91.13 2.88
DiT DRGG 78.46 84.94 6.03
DRGG-BE 85.55 87.58 4.13
DRGG-BS 82.08 88.65 5.08
DRGG-BEBS 88.84 90.82 3.20
InternImage DRGG 80.64 86.04 5.39
DRGG-BE 87.17 88.88 3.61
DRGG-BS 84.14 89.74 4.43
DRGG-BEBS 89.48 91.09 2.91
@@
ResNet-50 85.6 (+4.1) 94.3 (+0.4) 96.7 (+2.0) 92.8 (+2.3) 85.0 (+4.0) 76.7 (+0.0) 85.0 (+2.5) 79.8 (+5.5)
ViT 87.4 (+2.8) 95.2 (+0.3) 96.9 (+1.2) 93.3 (+2.1) 86.0 (+2.6) 78.3 (+3.3) 85.3 (+1.9) 81.2 (+6.1)
Swin 87.1 (+2.3) 95.0 (-0.4) 97.0 (+1.0) 92.8 (+0.9)⋆⋆ 86.5 (+2.3) 76.7 (-3.3) 86.6 (+1.4) 80.5 (+3.6)
DiT 85.0 (+2.6) 94.5 (+0.6) 96.5 (+1.5) 92.9 (+2.6) 84.4 (+3.0) 75.0 (-1.7) 85.3 (+3.3)⋆⋆ 78.8 (+5.0)
InternImage 87.1 (+2.3) 94.6 (-0.1) 96.6 (+0.8) 92.9 (+1.3) 86.9 (+2.8) 78.3 (+3.3) 86.4 (+2.2)⋆⋆ 79.7 (+3.1)
@@
ResNet-50 95.0 (+1.9) 83.6 (+4.1) 76.1 (+6.6) 84.6 (+5.0)⋆⋆ 84.8 (+5.1)⋆⋆ 87.4 (+2.4)
ViT 95.5 (+1.4) 85.1 (+3.2) 77.4 (+3.4)⋆⋆ 83.6 (+7.0) 86.2 (+3.6) 86.6 (+4.7)⋆⋆
Swin 95.5 (+1.0) 85.0 (+2.3) 77.1 (+1.9) 84.1 (+2.0) 85.5 (+0.7) 87.4 (+4.7)⋆⋆
DiT 94.6 (+1.5) 83.9 (+3.7) 75.9 (+4.9) 83.1 (+4.0)⋆⋆ 84.1 (+3.6) 89.0 (+0.8)
InternImage 95.0 (+1.0) 85.0 (+2.2) 77.4 (+3.2)⋆⋆ 89.1 (+3.0) 84.8 (+2.2) 85.0 (+3.1)
@@
ResNet-50 93.6 (+13.6) 95.2 (+8.9) 97.8 (+9.4) 93.3 (+13.2) 80.3 (+18.9) 90.9 (+18.2)⋆⋆ 87.6 (+13.5) 79.0 (+8.3)⋆⋆
ViT 93.8 (+5.1) 96.3 (+3.6) 98.3 (+5.5) 95.6 (+9.2) 82.7 (+5.5) 84.1 (+6.8) 87.6 (+19.5) 82.9 (+10.2)
Swin 94.1 (+7.3) 96.3 (+5.9) 98.2 (+6.6) 93.5 (+8.8) 77.2 (+7.9)⋆⋆ 86.4 (+13.6)⋆⋆ 89.7 (+21.1) 87.8 (+13.2)
DiT 92.6 (+7.8) 95.9 (+6.7) 97.6 (+7.2) 94.8 (+11.1) 85.8 (+15.7) 81.8 (+25.0) 87.0 (+30.3) 80.5 (+9.3)
InternImage 93.5 (+6.0) 96.2 (+5.6) 98.2 (+6.1) 95.4 (+10.7) 81.9 (+6.3) 86.4 (+13.6)⋆⋆ 88.6 (+13.5) 88.8 (+16.1)
@@
ResNet-50 97.9 (+6.4) 92.4 (+12.9) 91.9 (+15.5) 91.1 (+13.7) 90.4 (+17.1) 93.7 (+28.3)
ViT 98.4 (+4.9) 92.4 (+3.9) 93.8 (+6.7) 93.2 (+5.5) 92.9 (+9.6) 95.1 (+23.4)
Swin 98.6 (+5.8) 93.1 (+7.2) 93.4 (+8.6) 91.5 (+7.7) 92.2 (+12.7) 94.1 (+20.5)
DiT 97.9 (+7.2) 92.6 (+7.8) 91.6 (+9.1) 91.8 (+7.9) 91.9 (+15.2) 94.6 (+20.0)
InternImage 98.5 (+5.2) 92.9 (+6.8) 93.0 (+8.2) 92.8 (+7.7) 94.1 (+13.4) 95.1 (+17.6)
@@ 5.3.1 Analysis of Reading Order Prediction Improvements by Beam Search
Figure 7: Distribution of reading order prediction failures for each GT BBox direction. Failures from all backbones using DRGG-BEBS are accumulated in this visualization. These failures account for 7.09% of all predictions.
Figure 8: Distribution of reading order prediction failures for each GT BBox distance. These failures account for 7.09% of all predictions. The format follows Fig. 7.
@@ 5.3.2 Analysis of Parent-Child Prediction Improvements by BBox Embedding
Tables 56 show that accuracy for spatially challenging relations improved more than that for downward or short-distance relations6. These results indicate the effectiveness of BBox information in improving performance on such challenging cases.
@@
Figure 9: Distribution of parent-child prediction failures for each GT BBox direction. These failures account for 3.74% of all predictions. The format follows Fig. 7.
Figure 10: Distribution of parent-child prediction failures for each GT BBox distance. These failures account for 3.74% of all predictions. The format follows Fig. 7.
Figure 11: Distribution of parent-child prediction failures for each GT BBox category. These failures account for 3.74% of all predictions. The format follows Fig. 7.
@@ A.1 Annotation Guideline
A.2 Detailed Statistics of Dataset
@@
A.2.1 Layout Tree Statistics
@@
A.2.2 Reading Order Statistics
@@
A.2.3 Parent-Child Relation Statistics
@@
Right 2,143.42 889.82 41.14 5.22 0.38 0.00
Bottom-Right 2,045.98 151.06 23.44 7.77 4.46 0.76
Bottom 7,796.71 1,914.02 200.74 79.23 48.15 22.93
Bottom-Left 2,601.45 484.91 78.97 22.80 13.63 5.35
Left 1,565.02 521.84 247.10 57.44 4.59 0.38
Top-Left 131.19 13.88 8.79 7.13 4.71 1.15
Top 390.65 36.05 73.62 122.66 79.35 96.93
Top-Right 389.50 513.57 261.50 190.29 117.69 117.69
@@
Title 0.25 978.98 6.24 11.97 0.76 0.89 0.00 0.00
Author Info 0.00 0.76 820.53 125.59 7.39 23.31 0.89 0.00
Section 0.00 0.00 27.64 2,814.93 1,497.64 803.08 101.64 0.00
Text 0.00 0.00 1,951.85 1,488.22 538.02 1,990.70 283.28 0.00
List 0.00 0.00 996.69 426.82 439.18 647.31 101.39 0.00
Figure 0.00 0.00 826.26 1,273.72 364.54 767.04 65.60 1,555.98
Table 0.00 0.00 106.87 128.26 38.34 65.47 47.00 327.86
Caption 0.00 0.00 529.74 444.02 140.49 623.74 120.75 23.44
@@
Right 1,174.75 569.23 311.81 89.80 10.19 0.51
Bottom-Right 2,256.66 734.56 501.72 269.27 121.51 57.83
Bottom 3,367.21 1,016.05 1,142.78 850.59 494.59 329.38
Bottom-Left 564.39 198.96 129.54 70.95 50.06 25.35
Left 158.58 66.49 16.69 3.06 0.00 0.00
Top-Left 114.76 4.08 1.27 1.15 0.00 0.00
Top 369.63 6.50 2.55 4.71 2.42 2.04
Top-Right 43.05 140.62 85.85 42.03 15.92 15.03
@@
Root 999.11 980.00 5,266.08 901.67 206.09 672.27 75.79 0.00
Section 0.00 0.00 0.00 5,812.13 2,820.28 4,249.65 644.76 0.00
Figure 0.00 0.00 0.00 0.00 0.00 0.00 0.00 1,577.76
Table 0.00 0.00 0.00 0.00 0.00 0.00 0.00 329.51
@@ C.1 Effect of Text Embedding
ResNet-50 DRGG 68.75 75.05 8.82
DRGG-BE 84.46 86.61 4.31
DRGG-BS 76.83 83.12 6.66
DRGG-BEBS 88.45 90.38 3.20
DRGG-TE 71.19∗∗ 77.19∗∗ 8.20∗∗
DRGG-TEBS 78.49 84.29 6.14
DRGG-BETE 82.83 85.18 4.80
DRGG-BETEBS 86.83 88.87 3.69
ViT DRGG 80.41 85.84 5.41
DRGG-BE 86.24 88.11 3.86
DRGG-BS 83.78 89.25 4.49
DRGG-BEBS 89.67 91.34 2.86
DRGG-TE 80.16 86.27 5.57
DRGG-TEBS 83.13 89.05 4.75
DRGG-BETE 85.61 87.69 3.99
DRGG-BETEBS 88.66 90.51 3.11
Swin DRGG 79.84 85.78 5.70
DRGG-BE 87.14 88.86 3.57
DRGG-BS 83.11 89.02 4.75
DRGG-BEBS 89.50 91.13 2.88
DRGG-TE 80.20 85.70 5.52
DRGG-TEBS 84.09 89.50 4.44
DRGG-BETE 84.70 86.55 4.18
DRGG-BETEBS 88.28 90.06 3.26
DiT DRGG 78.46 84.94 6.03
DRGG-BE 85.55 87.58 4.13
DRGG-BS 82.08 88.65 5.08
DRGG-BEBS 88.84 90.82 3.20
DRGG-TE 79.17 85.43 5.82
DRGG-TEBS 82.06 88.47 5.10
DRGG-BETE 85.67 87.76 4.03
DRGG-BETEBS 88.67 90.60 3.20
InternImage DRGG 80.64 86.04 5.39
DRGG-BE 87.17 88.88 3.61
DRGG-BS 84.14 89.74 4.43
DRGG-BEBS 89.48 91.09 2.91
DRGG-TE 81.72 86.99 5.08
DRGG-TEBS 85.25 90.34 4.12
DRGG-BETE 84.85 86.82 4.21
DRGG-BETEBS 88.41 90.17 3.19
C.2 Additional Challenging Examples

### v2

# prompt[Prompt][List of Prompts]

@@ 3.2 Dataset Annotation
We construct SciPostLayoutTree by adding structural annotations to SciPostLayout [26]. In contrast to business or artistic posters, scientific posters have an explicit and consistent reading order intended by the authors. This property allows us to define a unique DFS-ordered tree structure for each poster. Each poster was annotated as a DFS-ordered tree rooted at the Root node [^1] See the Appendix for the annotation guideline.. The annotation was conducted by professional annotators from an external vendor, with one annotator assigned per poster. All annotations were reviewed by in-house supervisors and the authors with expertise in scientific poster layouts, ensuring rigorous quality control. To quantify annotation consistency, we evaluated 100 randomly selected posters by comparing the original annotations with independently annotated trees by two additional annotators, and observed a high agreement score (0.91) [^2] See the Appendix for the evaluation details.. Posters containing only a single BBox were excluded as structurally trivial. As a result, we obtained 7,849 annotated posters, split into 6,853/498/498 for training, validation, and test sets. We confirmed that 98% of posters in the validation and test sets have non-overlapping authors. The test set size of 498 posters is sufficient compared to prior datasets such as DocHieNet.
@@ 3.3.1 Layout Element Statistics
Title 7,842 1.00 ($\pm$ 0.04)
Author Info 7,692 0.98 ($\pm$ 0.15)
Section 41,321 5.26 ($\pm$ 2.43)
Text 52,725 6.72 ($\pm$ 4.89)
List 23,761 3.03 ($\pm$ 3.31)
Figure 38,625 4.92 ($\pm$ 3.62)
Table 5,655 0.72 ($\pm$ 1.26)
Caption 14,966 1.91 ($\pm$ 2.66)
All 192,587 24.54 ($\pm$ 9.58)
@@ 5.1 Experimental Settings
We compared four DRGG variants, each with a different combination of the BBox Embedding and beam search  [^5] We also evaluated VLMs (GPT-5, Gemini 3 Pro), but they showed lower performance than DRGG-based models. Details are provided in the Appendix..
@@ 5.2 Evaluation Metrics
ResNet-50 DRGG 68.74 75.07 8.83
DRGG-BE 84.24 86.44 4.41
DRGG-BS 76.79 83.14 6.65
DRGG-BEBS 88.45 90.40 3.22
ViT DRGG 80.40 85.94 5.45
DRGG-BE 86.38 88.24 3.85
DRGG-BS 83.89 89.46 4.46
DRGG-BEBS 90.04 91.73 2.78
Swin DRGG 79.90 85.77 5.69
DRGG-BE 86.73 88.42 3.69
DRGG-BS 83.11 89.02 4.74
DRGG-BEBS 89.26 90.88 2.95
DiT DRGG 78.33 84.81 6.09
DRGG-BE 85.32 87.36 4.22
DRGG-BS 82.12 88.72 5.07
DRGG-BEBS 88.69 90.63 3.24
InternImage DRGG 80.19 85.53 5.53
DRGG-BE 86.89 88.56 3.70
DRGG-BS 83.75 89.25 4.55
DRGG-BEBS 89.34 90.97 2.93
@@
ResNet-50 86.0 (+4.0) 94.5 (+0.6) 96.7 (+2.0) 92.7 (+2.4) 84.2 (+3.6) 73.3 (-1.7) 87.5 (+2.9) 79.8 (+5.5)
ViT 87.7 (+2.8) 95.2 (+0.0) 97.2 (+1.3) 93.1 (+2.0) 85.4 (+2.6) 75.0 (+5.0) 89.3 (+2.3) 81.4 (+5.7)
Swin 87.9 (+2.4) 95.0 (-0.5) 97.0 (+1.1) 92.7 (+1.0)⋆⋆ 85.5 (+2.8) 73.3 (-5.0) 89.0 (+2.0) 81.1 (+3.7)
DiT 85.3 (+2.8) 94.6 (+0.5) 96.6 (+1.6) 92.9 (+2.7) 83.3 (+3.3) 71.7 (-3.3) 88.1 (+2.3) 79.3 (+5.5)
InternImage 87.5 (+2.1) 94.8 (+0.0) 96.4 (+0.7) 92.9 (+1.3) 85.9 (+2.9) 73.3 (+3.3) 88.1 (+2.3)⋆⋆ 80.1 (+3.0)
@@
ResNet-50 94.7 (+1.9) 84.0 (+4.1) 77.4 (+6.4) 85.4 (+5.6)⋆⋆ 86.3 (+5.3)⋆⋆ 89.4 (+3.3)
ViT 95.3 (+1.3) 85.5 (+3.1) 80.5 (+4.4) 86.4 (+8.1) 87.0 (+2.3) 89.4 (+4.9)⋆⋆
Swin 95.2 (+1.0) 85.5 (+2.6) 78.7 (+2.2) 84.3 (+3.0) 87.0 (+1.5) 90.2 (+4.9)⋆⋆
DiT 94.4 (+1.6) 84.1 (+3.8) 77.2 (+5.1) 83.3 (+3.5) 87.8 (+3.8) 91.1 (+0.8)
InternImage 94.8 (+0.9) 85.1 (+2.1) 78.0 (+4.0) 88.9 (+2.5) 86.3 (+2.3) 87.8 (+3.3)
@@
ResNet-50 93.3 (+13.3) 95.3 (+8.4) 98.0 (+9.2) 92.6 (+13.6) 74.5 (+19.1) 90.7 (+18.6)⋆⋆ 88.8 (+14.0) 83.4 (+10.4)⋆⋆
ViT 94.3 (+5.3) 97.1 (+3.7) 98.4 (+5.5) 94.7 (+9.3) 77.3 (+7.1) 86.0 (+7.0) 89.9 (+20.8) 87.6 (+12.4)
Swin 94.8 (+7.9) 97.1 (+6.1) 98.5 (+6.6) 92.1 (+8.5) 73.0 (+9.9) 88.4 (+14.0)⋆⋆ 91.0 (+21.3) 90.2 (+14.0)
DiT 92.8 (+7.7) 96.3 (+6.5) 97.8 (+7.2) 94.7 (+12.1) 80.1 (+14.9) 83.7 (+23.3)⋆⋆ 90.4 (+32.6) 84.5 (+9.3)
InternImage 93.8 (+6.1) 96.9 (+5.5) 98.5 (+6.1) 94.3 (+10.8) 75.9 (+7.1)⋆⋆ 88.4 (+11.6) 90.4 (+13.5) 92.2 (+16.6)
@@
ResNet-50 97.8 (+6.4) 92.2 (+13.1) 92.0 (+14.7) 92.3 (+13.3) 91.1 (+16.5) 93.7 (+28.3)
ViT 98.3 (+5.1) 92.8 (+4.2) 94.8 (+6.6) 94.9 (+5.7) 94.0 (+9.5) 95.6 (+23.4)
Swin 98.5 (+5.8) 93.4 (+7.6) 94.2 (+9.0) 93.5 (+7.9) 93.4 (+12.3) 94.6 (+20.5)
DiT 97.9 (+7.2) 92.9 (+7.6) 92.4 (+9.5) 93.1 (+7.7) 93.0 (+15.2) 94.6 (+20.0)
InternImage 98.4 (+5.2) 93.1 (+7.1) 93.5 (+8.0) 94.6 (+7.5) 95.6 (+13.3) 95.6 (+17.6)
@@ 5.3.1 Analysis of Reading Order Prediction Improvements by Beam Search
Figure 7: Distribution of reading order prediction failures for each GT BBox direction. Failures from all backbones using DRGG-BEBS are accumulated in this visualization. These failures account for 7.00% of all predictions.
Figure 8: Distribution of reading order prediction failures for each GT BBox distance. These failures account for 7.00% of all predictions. The format follows Fig. 7.
@@ 5.3.2 Analysis of Parent-Child Prediction Improvements by BBox Embedding
Tables 56 show that accuracy for spatially challenging relations improved more than that for downward or short-distance relations8. These results indicate the effectiveness of BBox information in improving performance on such challenging cases.
@@
Figure 9: Distribution of parent-child prediction failures for each GT BBox direction. These failures account for 3.59% of all predictions. The format follows Fig. 7.
Figure 10: Distribution of parent-child prediction failures for each GT BBox distance. These failures account for 3.59% of all predictions. The format follows Fig. 7.
Figure 11: Distribution of parent-child prediction failures for each GT BBox category. These failures account for 3.59% of all predictions. The format follows Fig. 7.
@@ A.1 Annotation Guideline
A.2 Annotation Consistency
This subsection describes the evaluation procedure for the annotation consistency reported in Section 3.2. We evaluated agreement on 100 randomly selected posters from the test set by comparing the original annotations with independently annotated trees by two additional annotators. Agreement was computed for the entire tree, parent–child relations, and reading-order relations. Tree agreement was measured using STEDS defined in Section 5.2. Parent–child agreement was defined as the proportion of nodes whose assigned parent matches between two annotations. Reading-order agreement was defined as the proportion of consecutive node pairs in the DFS traversal that match between two annotations. Each score was computed per poster and averaged over the 100 posters. As a result, STEDS = 0.91, parent–child pair agreement = 0.97, and reading-order pair agreement = 0.94, indicating high agreement.
In addition, both the model STEDS in the evaluation experiments and the inter-annotator STEDS are approximately 0.9, raising the possibility that model errors may stem from inconsistent annotations. To investigate this hypothesis, we computed the Spearman rank correlation between per-poster model STEDS and inter-annotator STEDS. The average correlation is moderate (Spearman’s $\rho$ = 0.40), indicating that model errors are not primarily attributable to inconsistent annotations. The full correlation matrix is shown in Figure 12. As shown in the figure, correlations among models and among annotators are high, whereas correlations between models and annotators are low.
Figure 12: Heatmap of Spearman’s rank correlation coefficients between the STEDS of DRGG-BEBS models and those of the annotators.
A.3 Detailed Statistics of Dataset
@@
A.3.1 Layout Tree Statistics
@@
A.3.2 Reading Order Statistics
@@
A.3.3 Parent-Child Relation Statistics
@@
Right 2,143.97 889.16 41.15 5.48 0.38 0.00
Bottom-Right 2,045.61 150.72 23.57 7.64 4.46 0.76
Bottom 7,800.48 1,913.36 200.03 78.74 48.16 22.93
Bottom-Left 2,602.24 483.50 78.99 22.81 13.38 5.35
Left 1,568.61 524.53 247.29 57.20 4.59 0.38
Top-Left 130.97 14.27 8.28 7.01 4.20 1.15
Top 390.50 35.42 72.37 120.52 78.74 96.70
Top-Right 390.24 514.33 261.18 190.60 117.08 117.47
@@
Title 0.25 978.98 6.24 11.98 0.76 0.89 0.00 0.00
Author Info 0.00 0.76 820.61 125.62 7.39 23.19 0.89 0.00
Section 0.00 0.00 26.88 2,813.73 1,498.53 802.78 101.54 0.00
Text 0.00 0.00 1,946.24 1,491.02 538.03 1,991.97 283.48 0.00
List 0.00 0.00 1,000.89 427.83 439.55 646.83 101.67 0.00
Figure 0.00 0.00 826.22 1,274.30 364.12 766.85 65.23 1,555.61
Table 0.00 0.00 106.77 128.55 38.35 64.85 47.39 327.81
Caption 0.00 0.00 530.39 444.13 140.53 623.26 120.27 23.32
@@
Right 1,173.53 568.73 312.01 89.69 10.19 0.51
Bottom-Right 2,256.08 734.74 501.08 269.08 121.29 57.84
Bottom 3,367.94 1,016.94 1,144.86 851.83 495.35 329.72
Bottom-Left 565.17 199.77 130.34 70.96 50.07 25.35
Left 159.38 67.40 16.94 3.19 0.00 0.00
Top-Left 114.79 3.95 1.27 1.15 0.00 0.00
Top 369.35 6.24 2.42 4.71 2.42 2.04
Top-Right 42.81 140.02 85.36 41.92 15.93 15.03
@@
Root 999.11 980.00 5,264.49 902.41 206.14 669.51 75.42 0.00
Section 0.00 0.00 0.00 5,815.01 2,821.12 4,251.50 645.05 0.00
Figure 0.00 0.00 0.00 0.00 0.00 0.00 0.00 1,577.27
Table 0.00 0.00 0.00 0.00 0.00 0.00 0.00 329.47
@@ C.1 Effect of Text Embedding
ResNet-50 DRGG 68.74 75.07 8.83
DRGG-BE 84.24 86.44 4.41
DRGG-BS 76.79 83.14 6.65
DRGG-BEBS 88.45 90.40 3.22
DRGG-TE 71.21 77.23 8.21
DRGG-TEBS 78.46 84.19 6.15
DRGG-BETE 82.66 85.08 4.90
DRGG-BETEBS 87.15 89.20 3.59
ViT DRGG 80.40 85.94 5.45
DRGG-BE 86.38 88.24 3.85
DRGG-BS 83.89 89.46 4.46
DRGG-BEBS 90.04 91.73 2.78
DRGG-TE 80.32 86.50 5.56
DRGG-TEBS 83.35 89.33 4.73
DRGG-BETE 85.74 87.82 4.00
DRGG-BETEBS 89.04 90.93 3.06
Swin DRGG 79.90 85.77 5.69
DRGG-BE 86.73 88.42 3.69
DRGG-BS 83.11 89.02 4.74
DRGG-BEBS 89.26 90.88 2.95
DRGG-TE 80.31 85.77 5.52
DRGG-TEBS 84.18 89.51 4.46
DRGG-BETE 84.72 86.51 4.22
DRGG-BETEBS 88.37 90.11 3.23
DiT DRGG 78.33 84.81 6.09
DRGG-BE 85.32 87.36 4.22
DRGG-BS 82.12 88.72 5.07
DRGG-BEBS 88.69 90.63 3.24
DRGG-TE 79.25 85.52 5.83
DRGG-TEBS 82.02 88.40 5.13
DRGG-BETE 85.56 87.71 4.10
DRGG-BETEBS 88.81 90.82 3.16
InternImage DRGG 80.19 85.53 5.53
DRGG-BE 86.89 88.56 3.70
DRGG-BS 83.75 89.25 4.55
DRGG-BEBS 89.34 90.97 2.93
DRGG-TE 81.62∗∗ 86.74 5.15
DRGG-TEBS 85.05 90.02 4.21
DRGG-BETE 84.69 86.63 4.26
DRGG-BETEBS 88.19 89.94 3.24
C.2 VLM Results
We explored PromptTree, a model for decoding DFS-ordered trees using the latest VLMs such as GPT-5999https://openai.com/gpt-5/ and Gemini 3 Pro [^10] https://deepmind.google/models/gemini/pro/. Figure 15 provides an overview of PromptTree. The model takes as input the following components: the poster image; a visualization of the BBoxes rendered on a white canvas of the same size as the poster, where BBoxes are color-coded by category and assigned indices after sorting in y–x order; the metadata of each BBox (including position, category, and BBox index); and OCR-extracted text. In addition, two retrieved examples are provided as supplementary inputs. These examples are selected from the training set by ranking candidate examples in descending order of Intersection over Union (IoU), which is computed via Hungarian algorithm between the BBoxes of the target image and those of each candidate. Each retrieved example contains the same input information as the target image, along with the corresponding ground-truth reading order and parent–child relations. The prompt template provided to the VLM is shown in Prompt 16.
Figure 15: Overview of PromptTree.
Table 15: Comparison between DRGG variants and PromptTree, a VLM-based model.
Backbone Decoder STEDS ($\uparrow$) REDS ($\uparrow$) TED ($\downarrow$)
ResNet-50 DRGG 68.74 75.07 8.83
DRGG-BEBS 88.45 90.40 3.22
ViT DRGG 80.40 85.94 5.45
DRGG-BEBS 90.04 91.73 2.78
Swin DRGG 79.90 85.77 5.69
DRGG-BEBS 89.26 90.88 2.95
DiT DRGG 78.33 84.81 6.09
DRGG-BEBS 88.69 90.63 3.24
InternImage DRGG 80.19 85.53 5.53
DRGG-BEBS 89.34 90.97 2.93
GPT-5 76.96 83.13 6.34
Gemini 3 Pro 81.96 89.66 5.01
Table 15 presents a comparison between DRGG variants and PromptTree. PromptTree shows lower performance than all DRGG-BEBS variants when evaluated with either GPT-5 or Gemini 3 Pro, indicating that poster structure analysis remains challenging for the latest VLMs.
You are an expert in document structure analysis.
You are given bounding boxes from a scientific poster and one or more images of
the poster.
Your task is to infer:
1) the reading order of the bboxes,
2) the parent-child relationships forming a rooted tree.
Example 1:
Input bboxes:
- bbox_number=<bbox_id>, category=<category_name>, x=<x_norm>, y=<y_norm>,
w=<w_norm>, h=<h_norm>, text="<ocr_text_if_any>"
Images for Example 1:
[POSTER_IMAGE]
[BBOXES_IMAGE]
Output:
reading_order = <reading_order_list>
[…]
Example 2:
Input bboxes:
- bbox_number=<bbox_id>, category=<category_name>, x=<x_norm>, y=<y_norm>,
w=<w_norm>, h=<h_norm>, text="<ocr_text_if_any>"
Images for Example 2:
[POSTER_IMAGE]
[BBOXES_IMAGE]
Output:
reading_order = <reading_order_list>
[…]
Now solve the following example.
Input bboxes:
- bbox_number=<bbox_id>, category=<category_name>, x=<x_norm>, y=<y_norm>,
w=<w_norm>, h=<h_norm>, text="<ocr_text_if_any>"
Use both the textual description of the bboxes and the visual layout in the
images.
Output the result in the same format as the previous examples, using JSON only.
Do not include any explanations.
Images for the target example:
[POSTER_IMAGE]
[BBOXES_IMAGE]
{prompt}
Prompt used for decoding a DFS-ordered tree with PromptTree.
C.3 Additional Challenging Examples